OpenStudy (anonymous):

The volume (V) of a gas varies inversely as the pressure (P) upon it. The volume of a gas is 200 cubic centimeters under a pressure of 32 kg per square centimeter. What will be its volume under a pressure of 20 kg per square centimeter?

OpenStudy (shayaan_mustafa):

Hi Richiedguez :) Need help buddy?

OpenStudy (anonymous):

320 cm3

OpenStudy (shayaan_mustafa):

\[\large V \alpha \frac{1}{P}-------(1)\]When proportional sign is changed to = then proportionality constant is used. Let it is k\[\large V=k\frac{1}{P}\]Find "k" by substituting V=200cm3 and P=32kg/cm2\[\large k=6400\]Now again use eq (1) to find V by substituting k=6400 and P=20kg/cm2 Good luck :)